# Customizable wave tailoring materials enabled by coupling nonlinear inverse design at two scales

### v1

@@ 2.3 Shape optimization for desired nonlinear constitutive law
In the example shown herein, we strive for generality, thus there was no single impact condition specified. The solutions were chosen which exhibited low impactor mass and velocity, as well as low maximum strain, in order to avoid the onset of plasticity in the fabricated structures. One can just as well begin with a strict impact condition requirement, and flow through the multilevel optimization methodology accordingly. If larger mass and velocity, and therefore larger strains, are desired, the mesostructure optimization portion can be undertaken with a Neo-Hookean material model, and the real-world structures fabricated out of some soft, flexible material, such as silicone (see Ref. [23]).
2.4 Experimental validation
In our experimental realization of the chains, the spring scale $a_{s}=40$ mm was chosen and incorporated into the rigid frame and connector, resulting in a unit cell length $a=125$ mm. In order to minimize the non-longitudinal motion of the masses, a second nonlinear spring was added in parallel to the first. A single unit cell (shown in Fig. 4C) was fabricated by cutting the shape out of a polycarbonate sheet via computer-numerical-control (CNC) milling. We note the design of the frame surrounding a spring is particularly important in properly imposing boundary conditions, and thus critical in matching simulation predicted force displacement curves. This single unit cell was tested quasi-statically to confirm the performance of the spring, as shown in Fig. 4D. Repeated, cyclic quasi-static loading tests were performed to confirm that no onset of plasticity or fatigue occurs during the dynamic experiment (see SI Note 5).
@@ 2.4 Experimental validation
A chain of twenty unit cells was fabricated and hung from a frame in order to minimize friction, as shown in Fig. 4E. Stationary rails (not pictured in Fig. 4E) were positioned on either side of the hanging masses to minimize lateral motion, with care taken to position the rails such that contact with the hanging masses was nonexistent at rest. Impact tests were undertaken where the mass and velocity of the impactor were chosen based upon the optimal conditions identified in the DEM results, in the case with an impactor mass of $M=20$ g and a velocity of $V=2$ m/s. Data was collected through the use of several cameras positioned along the length of the chain, allowing digital image processing to be used to track the impact wave across the length of the system. In addition, a laser Doppler vibrometer (see SI Note 6) was pointed at the last unit cell in the system, allowing for the second measurement of the velocity of the last unit cell. Impact tests were repeated several times. A similar chain of twenty linear unit cells, with similar linear stiffness and mass (the mass of the linear and nonlinear unit cells are $152.5$ g and $152.8$ g, respectively) values, was then constructed to act as a control for comparison against the nonlinear chain (see SI Note 6 for more details), and the impact tests were repeated. It is noted that for the kinetic energy transmission ratio considered herein, the magnitude of the linear stiffness (even if different between the linear and nonlinear chain) does not matter (see SI Note 7).
@@
Figure 5: Experimental validation of the optimal nonlinear and linear chains, compared with simulation predictions. A-D) Spatiotemporal evolution of kinetic energy in the system (normalized by the input kinetic energy, or initial total energy). The nonlinear chain is shown in A (experiment, nonlinear trial 1) and B (simulation), while the linear case is shown in C (experiment, linear trial 1) and D (simulation). The experimental spatiotemporal plots include a smoothing of displacement values to assist with noise induced by differentiating the discrete time camera data. The spatiotemporal plots for other trials are included in SI Note 6. E) The maximum kinetic energy (linear/nonlinear) seen at each unit cell for both experiment and simulation. A value greater than 1 indicates superior performance of the nonlinear chain as compared to the linear. The experiment values are the average taken from three experimental trials. The X marks the average experimental value recorded by the vibrometer (see SI Note 6 for the full data sets), which collected data from only the last unit cell. F) The simulated sensitivity of the $KE$ ratio (truncated at 1) to impact conditions, wherein the simulation was run with the coefficients that were found experimentally from the nonlinear spring in Fig. 4D (that is, the physically achieved coefficients).
@@
Several key metrics were examined to confirm the performance of the system, the results of which are summarized in Fig.  5. Foremost among these results is the velocity (or kinetic energy) which was transmitted to the end of the chain. Measured spatiotemporal kinetic energy responses are shown in Fig. 5A-D, in which we can see that the nonlinear cases dissipate and trap kinetic energy through unit cell snapping, preventing much of it from reaching the right (or protected) end of the chain (a large portion of the kinetic energy is seen to remain, reflecting back and forth, in the first 4-5 unit cells in panels A and B). We note an excellent match between simulation (using the experimentally fit coefficients taken from the quasi-static force-displacement curve shown in Fig. 4D) and experiment in these spatiotemporal plots. In particular, as seen in Fig. 5A and B, the number of unit cells that underwent snap-through in the nonlinear chain (four in both simulation and experiment, identified by the high kinetic energy values in unit cells 1-4 which then greatly reduce in subsequent unit cells when snapping is no longer occurring—see also the strain x-t diagrams and the experiment videos included in SI Note 6) also shows excellent agreement. Computational explorations showed this metric to be highly sensitive to small changes in polynomial coefficients and input conditions. Figure 5E shows a comparison of the ratio (of linear to nonlinear cases) of the maximum kinetic energy seen at each unit cell, in both experiment (averaged across three trials—we note that several later trials [See SI Note 6] were excluded from averaging, due to a likely degraded behavior of the spring after multiple tests, slightly adjusted experimental conditions, and potential sensitivity effects as explored further below) and simulation. We observe largely good agreement between the simulation and experiment. We see a larger discrepancy at the last particle (Fig. 5E), which we attribute to non-ideal boundary conditions. A point of particular note, is that although the targeted conditions show excellent predicted performance, the behavior can be sensitive to small variations in impactor mass and velocity (see the simulation data of Fig. 5F). To more quantitatively describe this sensitivity, in SI Note 3, we calculate gradient of Fig. 5F, which shows $|\partial(KE\text{ratio})/\partial M|\approx 2$ and $\partial(KE\text{ratio})/\partial V\approx 5$. This sensitivity is amplified in the presence of bistability, in which the snapping (or lack thereof) of a single unit cell can push the system from one qualitative behavior well to another (see SI Note 3). The impactor velocities seen herein were not precise (ranging from $2.10-2.19$ m/s, see SI Note 6), and this, coupled with the aforementioned sensitivity provides an insight into variations in chain performance—namely, that the slight variations in impactor velocity we see in experiment have the potential to easily knock the system out of its optimal performance region, into one in which poorer performance is to be expected. Despite the sensitivity of the system, per Fig. 5E the kinetic energy ratio ($KE_{lin}/KE_{non})$ remains greater than 1 (superior performance of the nonlinear chain compared to the linear) for both simulation and experiment once a critical number of unit cells has passed (unit cell 5 for the simulation, unit cell 4 in experiment). Additional tests of the decay rate of a low amplitude propagating pulse in both chains (see SI Note 5) suggest the damping in the linear chain is actually higher than the nonlinear one. This highlights the value of nonlinear wave manipulation, where superior performance can be seen even compared to a linear system with greater damping.
@@
We include here a further discussion on the simulated sensitivity of the performance with respect to the impact conditions plotted in Fig. 5F (the gradients of $KE$ ratio with respect to the mass and velocity are represented in SI Note 3). The variability of the mass-velocity space is immediately apparent, with several very small regions of excellent performance (low $KE$ ratios) surrounded by oscillating regions of lower performance (relatively higher $KE$ ratios), and even several points of poor performance (in which the ratio of nonlinear to linear maximum kinetic energy approaches or exceeds 1). The optimum performance region targeted in this work lies in one of these small low $KE$ ratio regions, and slight deviations in the impactor velocity and mass can therefore be expected to produce variable and poorer nonlinear chain performance. This phenomenon reveals important characteristics regarding the sensitivity of the system, and more generally nonlinear dynamical systems wherein bifurcation can cause sharp changes in behavior [38]. We believe the proximity of regions of poor performance to regions of good performance (e.g. Fig. 5F) motivates a consideration of nearby conditions in future work. For instance, we expect there are application scenarios in which a region of reduced sensitivity to changes in stimuli may be desirable at the expense of slightly lowered performance.
3 Conclusion
@@
In this work, we have demonstrated an integrated, two-level inverse design methodology for nonlinear wave tailoring in a mechanical system. Including the possibility of more highly discretized materials ($100$ unit cells, see SI Note 4), our optimizations of nonlinear constitutive laws suggest the potential for over two orders of magnitude improvement (reduction) in kinetic energy transmission via the use of a nonlinear material compared to a comparative linear material. Our initial experiments on materials with fewer unit cells ($20$) show close agreement with our simulations, validating our two-level design approach. We also observed large sensitivity of the kinetic energy transmission metric to small changes in nonlinearity and impact conditions, which we suggest highlights the need for high precision, free-form nonlinearity tailoring, as well as future study of nonlinear dynamical design for robustness-of-response to the presence of defects, disorder, and stimuli variation. Within the considered topic of impact mitigation, a near term question of future interest would be how such optimal nonlinearities would change for a different metric such as maximum transmitted force or peak tensile stress anywhere in the material. Similarly, although our proof of concept study focused on impact, we envision a wide range of applications that could make use of precisely tailored nonlinear signal transformation such as mechanical computing, acoustic signal and image processing, vibration mitigation, and materials that autonomously respond and conduct directed work via rapid shape-change. Additional future research areas include extension of this method to less scale separated dynamical (no lumped masses), higher-dimensional, irreversible constitutive, heterogeneous material, active, and coupled-physics settings.
@@ Acknowledgements
Conflict of Interest
The authors declare no conflicts of interest.
[…]

### v2

@@ Results
In the example shown herein, we strive for generality, thus there was no single impact condition specified. The solutions were chosen which exhibited low impactor mass and velocity, relatively smooth sensitivity to impact conditions (discussed in more detail below), and low maximum strain in order to avoid the onset of plasticity in the fabricated structures. One can just as well begin with a strict impact condition requirement, and flow through the multilevel optimization methodology accordingly. If larger mass and velocity, and therefore larger strains, are desired, the mesostructure optimization portion can be undertaken with a Neo-Hookean material model, and the real-world structures fabricated out of some soft, flexible material, such as silicone (see Ref. [23]).
Experimental validation
@@
In our experimental realization of the chains, the spring scale $a_{s}=59$ mm was chosen and incorporated into the rigid frame and connector, resulting in a unit cell length $a=125$ mm. A chain of twenty unit cells (a single unit cell is shown in Fig. 4C) was fabricated and hung from a frame in order to minimize friction, as shown in Fig. 4E (see Note 7 in the Supplementary Information for a more detailed breakdown of the chain design). Impact tests were undertaken where the mass and velocity of the impactor were chosen based upon the optimal conditions identified in the DEM results, in the case with an impactor mass of $M=40$ g and a velocity of $V=1.37$ m/s. Data was collected through the use of several cameras positioned along the length of the chain, allowing digital image processing to be used to track the impact wave across the length of the system. In addition, a laser Doppler vibrometer (see Supplementary Information Note 7) was pointed at the last unit cell in the system, allowing for a second measurement of the velocity of the last unit cell. Impact tests were repeated several times. A similar chain of twenty linear unit cells, with similar linear stiffness and mass (the mass of the linear and nonlinear unit cells are $400.4$ g and $398.8$ g, respectively) values, was then constructed to act as a control for comparison against the nonlinear chain (see Supplementary Information Note 6 for more details), and the impact tests were repeated. It is noted that for the kinetic energy transmission ratio considered herein, the magnitude of the linear stiffness (even if different between the linear and nonlinear chain) does not matter (see Supplementary Information Note 8).
@@
Figure 5: Experimental validation of the optimal nonlinear and linear chains, compared with simulation. A-D) Spatiotemporal evolution of kinetic energy in the system (normalized by the input kinetic energy, or initial total energy). The nonlinear chain is shown in A (experiment, nonlinear trial 1) and B (simulation), while the linear case is shown in C (experiment, linear trial 1) and D (simulation). The experimental spatiotemporal plots include a smoothing of displacement values to assist with noise induced by differentiating the discrete time camera data. The spatiotemporal plots for other trials are included in Supplementary Information Note 7. E) A ratio of maximum kinetic energy (linear/nonlinear) seen at each unit cell for both experiment and simulation. A value greater than 1 indicates superior performance of the nonlinear chain as compared to the linear. The experiment values are the average taken from three experimental trials. The X marks the average experimental value recorded by the vibrometer (see Supplementary Information Note 7 for the full data sets), which collected data from only the last unit cell. F) The simulated sensitivity of the $KE$ ratio (truncated at 1) to impact conditions, wherein the simulation was run with the coefficients that were found experimentally from the nonlinear spring in Fig. 4D (that is, the physically achieved coefficients).
@@
Several key metrics were examined to confirm the performance of the system, the results of which are summarized in Fig.  5. Foremost among these results is the velocity (or kinetic energy) which was transmitted to the end of the chain. Measured spatiotemporal kinetic energy responses are shown in Fig. 5A-D, in which we can see that the nonlinear cases dissipate and trap kinetic energy through unit cell snapping, preventing much of it from reaching the right (or protected) end of the chain (a large portion of the kinetic energy is seen to remain, reflecting back and forth, in the first 9-11 unit cells in panels A and B). We note an excellent match between simulation (using the experimentally fit coefficients taken from the quasi-static force-displacement curve shown in Fig. 4D) and experiment in these spatiotemporal plots. Figure 5E shows a comparison of the ratio (of linear to nonlinear cases) of the maximum kinetic energy seen at each unit cell, in both experiment (averaged across three trials) and simulation. We note that the ratio shown in Fig. 5E is the inverse of $KE$ ratio, for ease of visualization, such that larger numbers denote better impact protection. We see a larger discrepancy at the last two particles (Fig. 5E), which we attribute to non-ideal boundary conditions. We note that damping characterization (see Methods section) suggest the damping in both chains is similar ($\zeta$ of 0.005 for nonlinear and 0.003 for linear). Further, we show in Supplementary Information Note 6, that were the linear chain to have higher damping, e.g., $\zeta=0.006$, this would have negligible effect on the $KE$ ratio. This highlights the value of nonlinear wave manipulation, where superior performance can be seen without heavy reliance upon damping.
@@
A point of particular note, is that although the targeted conditions show excellent predicted performance, the behavior can be sensitive to small variations in impactor mass and velocity, as seen in the simulation data of Fig. 5F. The variability of the mass-velocity space is immediately apparent, with several very small regions of excellent performance (low $KE$ ratios) surrounded by oscillating regions of lower performance (relatively higher $KE$ ratios), and even several points of poor performance ($KE$ ratio $>1$). The optimum performance region targeted in this work sought out a region with relatively low drops in performance (relative to other regions explored via simulation), but nonetheless there still exist oscillatory performance impact conditions nearby. In order to more quantitatively describe this sensitivity, in Supplementary Information Note 4, we calculate the gradient of Fig. 5F, which shows $|\partial(\log_{10}(KE\text{ratio}))/\partial M|$ can reach near $1$ g-1 and $|\partial(\log_{10}(KE\text{ratio}))/\partial V|$ can reach up to $40$ s/m. This means that, for the most sensitive regions of the impact conditions landscape, a change in 1 g of impactor mass can result in an up to 6.7$\sim$ 6.7$\times$ 6.7 × change in $KE$ ratio, or a 0.01 m/s change in impactor velocity can result in an up to 2.5$\sim$ 2.5$\times$ 2.5 × change in $KE$ ratio. However, as with the case of the sensitivity to the nonlinear stiffness parameters, near our chosen optimum, the sensitivity is significantly lower, with $\partial(\log_{10}(KE\text{ratio}))/\partial M=0.16$ g-1 and $\partial(\log_{10}(KE\text{ratio}))/\partial V=-7.8$ s/m. This sensitivity is amplified due to the presence of bistability, in which the snapping (or lack thereof) of a single unit cell can push the system from one qualitative behavior well to another (see Supplementary Information Notes 4 and 9 for more in depth analysis and an example, respectively). The impactor velocities seen herein were not precise (ranging from $1.36-1.41$ m/s, see Supplementary Information Note 7), and this, coupled with the aforementioned sensitivity provides an insight into variations in chain performance—namely, that the slight variations in impactor velocity we see in experiment have the potential to easily knock the system out of its optimal performance region, into one in which poorer performance is to be expected. This phenomenon reveals important characteristics regarding the sensitivity of the system, and more generally nonlinear dynamical systems wherein bifurcation can cause sharp changes in behavior [40]. We believe the proximity of regions of poor performance to regions of good performance (e.g., Fig. 5F) motivates a consideration of nearby conditions in future work. For instance, we expect there are application scenarios in which a region of reduced sensitivity to changes in stimuli may be desirable at the expense of slightly lowered performance. Despite the sensitivity of the system, however, per Fig. 5E the kinetic energy ratio ($KE_{lin}/KE_{non})$ remains greater than 1 (superior performance of the nonlinear chain compared to the linear) for both simulation and experiment once a critical number of unit cells has passed (unit cell 11 for the simulation, unit cell 10 in experiment).
Discussion
@@
In this work, we have demonstrated a bilevel inverse design methodology for nonlinear wave tailoring in a mechanical system. Including the possibility of more highly discretized materials (e.g., $100$ unit cells, see Supplementary Information Note 5), our optimizations of nonlinear constitutive laws suggest the potential for over two orders of magnitude improvement (reduction) in kinetic energy transmission via the use of a nonlinear material compared to a comparative linear material. Our initial experiments on materials with fewer unit cells ($20$) show close agreement with our simulations, validating our bilevel design approach. We also observed large sensitivity of the kinetic energy transmission metric to small changes in nonlinearity and impact conditions, which we suggest highlights the need for high precision, free-form nonlinearity tailoring, as well as future study of nonlinear dynamical design for robustness-of-response to the presence of defects, disorder, and stimuli variation. Within the considered topic of impact mitigation, a near term question of future interest would be how such optimal nonlinearities would change for a different metric such as maximum transmitted force or peak tensile stress anywhere in the material. Similarly, although our proof of concept study focused on impact, we envision a wide range of applications that could make use of precisely tailored nonlinear signal transformation such as mechanical computing, acoustic signal, and image processing, vibration mitigation, and materials that autonomously respond and conduct directed work via rapid shape-change. Additional future research areas include extension of this method to less scale separated dynamical (no lumped masses), higher-dimensional, irreversible constitutive, heterogeneous material, active, and coupled-physics settings.
Methods
DEM Simulations
We numerically integrate nondimensional Eq. (S8) in Supplementary Information Note 2 using a Runge-Kutta algorithm (ode45 in MATLAB) with impactor velocity $V/V_{0}$ and mass $M/m$ applied to the impactor particle. The output and maximum integration timestep is selected by estimating the highest nondimensional frequency $f_{max}=\frac{1}{2\pi}\sqrt{\frac{k_{max}}{M_{0}}}$, where $k_{max}=\max(1,1+2c_{2}+3c_{3})$, such that the timestep $\Delta t=1/(1000f_{max})$. The nondimensional displacement and velocity tolerances are set to $10^{-10}$. The total energy conservation is checked for an undamped DEM system, and deviation is less than 0.1% over the entire simulation duration.
Structural Optimization
We include here further details about the shape optimization of the spring geometry, while noting that a full description of the method can be found in our previous work [23]. A level-set method is used to perform a nonlinear, displacement control, 2D continuum shape optimization. The spring is modeled using a 2D plane stress condition, with a Kirchhoff material model (linear elastic material model, with geometric nonlinearity included), as we have found this material model to be a good representation of polycarbonate throughout displacement ranges which do not reach the plasticity threshold of the material. The optimization performed herein was solved with a uniform mesh consisting of $750\times 750$ quadrilateral elements. A symmetry boundary condition (rollers) is applied along the bottom edge of the optimization domain, such that only one half of the spring need be simulated. The top edge of the domain has a fixed boundary condition enforced. A 1D applied displacement is applied at the bottom right edge of the domain, where the connector to the adjacent spring would exist in the physical system. These conditions, along with the initial condition supplied to the optimizer, are illustrated in Note 6 in the Supplementary Information.
Prior to manufacture and experimental tests, the force displacement output from the optimization process was further confirmed via a FEM simulation in COMSOL Multiphysics. Following the verification process outlined in our previous work [23], the optimal level set was converted to a 2D geometry in COMSOL, and a quasi-static simulation was performed. This simulation similarly used 2D plane stress, with a Kirchhoff material model and the same applied boundary conditions as the optimization process. The COSMOL simulation, however, employed a body-fitted mesh consisting of 20,447 triangular elements. The results, shown in Fig. 4D, confirm the accuracy of the optimization simulation.
As outlined in Eq. 2, the objective function for the optimization is taken to be the ratio of polynomial coefficients resulting from the force-displacement curve of the current design. This polynomial is fit to the force-displacement curve using the equation $c=V^{-1}C$, in which $V$ is the Vandermonde matrix, and $C$ is compliance (we note that herein, we use lowercase $c$ to refer to polynomial coefficients, and uppercase $C$ to refer to compliance). Fitting to compliance, rather than force, is done for ease of sensitivity calculation, as compliance optimization is a well known optimization problem. Noting that $C=F\Delta x$ (in which $\Delta x$ is displacement), it is a simple matter to transition between force and compliance. We can then write the sensitivity for the polynomial ratio objective as a combination of known compliance sensitivity terms and sensitivity terms related to the fitting of the polynomial.
$\frac{\partial c_{i}}{\partial\Omega}=\sum_{j=1}^{m}\frac{\partial c_{i}}{\partial C_{j}}\frac{\partial C_{j}}{\partial\Omega}=\sum_{j=1}^{m}V^{-1}(i,j)\frac{\partial C_{j}}{\partial\Omega},$ (3)
where $\Omega$ refers to the level set domain, $i$ refers to the degree of polynomial, $j$ refers to the displacement step, and $m$ is the total number of displacement points under consideration (see our previous work for more details[23]).
Unit Cell Design and Manufacturing
After the nonlinear spring had been designed via optimization, the design was resized to the experiment scale (wherein the length of one spring, $a_{s}$, was taken to be 59 mm). The single spring design was reflected over the symmetry condition, and in order to minimize the non-longitudinal motion of the masses, a second set of nonlinear springs was added in parallel to the first. A single unit cell (shown in Fig. 4C) was fabricated by cutting the shape out of a polycarbonate sheet via computer-numerical-control (CNC) milling. This milled shape was then sanded down at the edges to fit smoothly into a rectangular stainless steel frame, which was commercially manufactured via sheet cutting, and secured in-plane with an acrylic backing bolted into the frame and glued to the milled spring (see Note 7 in the Supplementary Information for a visualization). We note the design of the frame surrounding a spring is particularly important in properly imposing boundary conditions, and thus critical in matching simulation predicted force displacement curves. The choice of a stainless steel frame herein was made in order to meet these criteria, while still allowing for the milling of the spring to fit within the manufacturing constraints of the manufacturing equipment.
A single unit cell was tested quasi-statically to confirm the performance of the spring, as shown in Fig. 4D. Repeated, cyclic quasi-static loading tests were also performed to confirm that no onset of plasticity or fatigue occurs during the dynamic experiment (see Supplementary Information Note 6).
Data Collection
The motion of each unit cell was captured with the use of cameras mounted above the chain. The length of the chain necessitated the use of multiple cameras to capture the motion all unit cells. Each experimental trial therefore has four videos associated with it, shot in slow motion at 240 fps on four iPhones, with each video overlapping by at least one unit cell to ensure proper spatial synchronization and the capture of all motion. For each trial, the cameras and unit cells were spaced such that camera 1 captured unit cells 1-6, camera 2 captured unit cells 4-11, camera 3 captured unit cells 9-16, and camera 4 captured unit cells 14-20. Video processing was performed in MATLAB. Temporal synchronization between each video was achieved by a visual signal which caused a fluctuation in intensity of light, which could then be detected in each video and used to synchronize the times between the recording devices. Tracking of each particle was achieved with the use of colored markers on each unit cell. Post processing in MATLAB allowed each frame to be separated into discrete color channels, allowing for the isolation and tracking of red, blue, and green markers (blue was used for the impactor, green for the back edge of each unit cell, and an additional red marker for unit cells which were covered by overlapping camera fields of view, in order to enable spatial synchronization of the data from each camera).
Damping Characterization
[…]
$L(f)=\frac{A}{\pi}\frac{\Gamma/2}{(f-f_{0})^{2}+(\Gamma/2)^{2}},$ (4)
where $A$ is the amplitude parameter, $f_{0}$ is the central frequency, and $\Gamma$ is the full width at half maximum in frequency. The fitting process was done by using the built-in fit command in MATLAB. The $Q$ factor was found by taking $f_{0}/\Gamma$ from the numerical fitting, which are approximately $92$ for the nonlinear and $177$ for the linear spring unit cell. Based on these $Q$ factors, by using the equation $\zeta=1/(2Q)$, we calculated the damping ratios $\zeta$ as $0.005$ and $0.003$ for nonlinear and linear chains, respectively, which were used in the numerical simulations.
@@ Acknowledgements
Author contributions statement
N.B., and H.A.K. designed research; B.M., H.X. I.F., K.Q. N.B., and H.A.K. performed research, K.Q., B.M., and I.F. conducted the experiments, B.M., H.X. I.F., K.Q. analysed the results; and B.M., H.X. I.F., K.Q. N.B., and H.A.K. wrote the paper.
Data Availability
Two sets of four videos (reduced resolution and sped up via downsampling by 8×8$\times$8 × to meet file size limitations) corresponding to the data shown in Fig. 5A and Fig. 5C are included as part of the Supplementary Information. The same naming convention is used as in Supplementary Information Table 1, but with the suffix “_SpedUpLowRes”.
All related videos regarding the data in the main text are uploaded to: Boechler, Nicholas (2024), “Customizable wave tailoring materials enabled by nonlinear bilevel inverse design 1”, Mendeley Data, V1, doi: 10.17632/2wgwfy2wfg.1
Videos related to the data in Supplementary Information Note 9 are uploaded to: Boechler, Nicholas (2024), “Customizable wave tailoring materials enabled by nonlinear bilevel inverse design 2”, Mendeley Data, V2, doi: 10.17632/6bg6hr5kyr.2
Competing Interests
The authors declare no competing interests.